OpenStudy (anonymous):

What is the sum of a 17-term arithmetic sequence where the first term is 7 and the last term is -105?

OpenStudy (anonymous):

formula is \[s _{n}=\frac{n}{2}(a _{1}+a _{n})\]

OpenStudy (anonymous):

use n=17 a1=7 an=-105

OpenStudy (anonymous):

let me know what you got ?

OpenStudy (anonymous):

\[\Large s _{17}=\frac{17}{2}(7+(-105))\]

OpenStudy (anonymous):

uhhh.... okay. so, n=12. s12 = 12/2 (or 6) times (7-105) ?

OpenStudy (anonymous):

oh wait. n is 17??

OpenStudy (anonymous):

dyslexia.. i could have sworn that said 12 :( so, 17/2 = 8.5 times -98?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-833 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

THANKS :D

OpenStudy (anonymous):

yw:)